SPECTRA OF ETHANE IN He DROPLETS IN THE 3 μm RANGE

\underline{LUIS GOMEZ}, EVGENIY LOGINOV, DMITRY SKVORTSOV, HIROMICHI HOSHINA, and AN-DREY F. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90089-0482.

The infrared spectra of ethane molecules embedded in He droplets have been studied via depletion spectroscopy in the ν = 2880-3000 cm^{−1} spectral range. Four features observed in the spectra are assigned as the ν_{5}, ν_{8+11}(perpendicular and parallel components), and ν_{7} vibrational bands. Band origins and rotational constants for ethane in He are obtained and compared with corresponding gas phase values. Spectra of large ethane clusters (C_{2}H_{6})_{n} (with n = 1300-13000) in He are also reported.